# Coward: Toward Practical Proactive Federated Backdoor Defense via Collision-based Watermark

### v1

@@ IV-C Watermark Interaction and Detection
V Experiments
V-A Main Settings
Datasets and Models. We conduct experiments on 3 classical image classification benchmark datasets, including EMNIST [58], CIFAR-10 and CIFAR-100 [59] . We employ ResNet-18 [60] as the default architecture.
@@ V-A Main Settings
Federated Training Configurations. All experiments adopt the standard FedAvg setting [1], where we simulate 100 clients in total, with 10 randomly selected to participate in each training round. To simulate the inherent data heterogeneity across clients, we follow [61] to use Dirichlet sampling to partition the dataset. The Dirichlet concentration parameter $\alpha$ controls the label distribution skewness, where smaller values indicate more unbalanced distributions. Specifically, we use $\alpha=0.9$ by default to simulate mild skewness, and decrease it to 0.3 to model more severe distribution shifts. During local training, each benign client performs 2 epochs with a learning rate of 0.03, while the global training proceeds for 1200 rounds. For conciseness, we omit the less critical parameter details and follow [18] for all settings not explicitly specified.
@@
Attack Configurations. To ensure a comprehensive evaluation, we consider both single-attacker and multi-attacker scenarios, each covering various attack strategies and backdoor types. (1) In the single-attacker scenario, we adopt a range of attack baselines, including the naive Vanilla training [23], the stealthy PGD attack [33], and the latest state-of-the-art methods Neurotoxin [34] and Chameleon [11]. Regarding the backdoor type, we adopt three diverse types of triggers, ranging from the traditional visible trigger (BadNet [23]), to the invisible trigger (Blend [62]), and the natural feature-based trigger (Semantic [63]), with increasing levels of imperceptibility. (2) In the multi-attacker scenario, we increase the ratio of attackers from 30% to 70%. To ensure comprehensive coverage, we include both aligned and divergent attack goals: Uniform [48] with homogeneous attackers, DBA [36] with a shared target but distributed triggers, and NBA [38] with non-cooperative attackers using distinct backdoors. Across all settings, poisoning uses a consistent learning rate of 0.03 by default, aligned with benign training, and reduces it to half for more challenging stealthy evaluation purposes. The backdoor target is class 0 by default. Similar to BackdoorIndicator, the malicious client launches the attack at the 1000th global round, with poisoning action lasting for 200 global rounds.
@@
Defense Baselines. We evaluate six representative defense methods that span a range of detection strategies, from classical model-space analysis to more advanced and proactive techniques: We begin with traditional methods that rely on the statistical properties of model updates: (1) MultiKrum [47] detects attackers based on the Euclidean distance among client updates. (2) Foolsgold [48] identifies malicious clients by analyzing pairwise cosine similarity to detect overly aligned update directions. (3) Rflbat [51] projects model updates into a PCA space and flags outliers as potential attackers. Then we further include more advanced approaches that incorporaFte prediction behavior or mitigation mechanisms: (4) Deepsight [16] evaluates similarity in model predictions and output neuron activations to reveal abnormal client behavior. (5) Flame [17] combines detection based on model similarity together with mitigation techniques such as gradient clipping and noise injection. Finally, we include the pioneering proactive method that shifts the detection paradigm: (6) BackdoorIndicator [18] injects OOD patterns into the global model and identifies attackers based on their reactions to these patterns.
@@
Our Configurations. We detail the key factors involved in our watermark injection and detection process. To construct the OOD planting set, we use EMNIST (expanded to three channels) when the main task is CIFAR-10 or CIFAR-100, and grayscale CIFAR-10 when the main task is EMNIST. The same planting set is used for BackdoorIndicator to ensure a fair comparison. By default, the planting set contains 1,000 samples. At each global round, the server injects the watermark by applying a trigger to 20% of the planting set, using a learning rate of 0.001 for 5 iterations. In our main experiments, we adopt WaNet [24] as the default watermark trigger. For detection, a fixed 5% threshold is used across all settings: clients with watermark accuracy below this are identified as attackers. The watermark target is set to 8 by default.
@@
V-B Main Results
For a comprehensive evaluation, We evaluate all baseline defenses from following three perspectives:
TABLE II: Detection performance against vanilla single-client attacks under different non-i.i.d. settings. Boldface indicates the best result value (excluding the collapsed case with TPR/FPR=0/0), while red highlights metric-specific failures based on a 50% threshold ($i.e.$, values above 50% for FPR and ASR, and below 50% for TPR and BA).
@@ V-B Main Results
TABLE III: Detection performance against stealthy single-client attacks with advanced training strategies. Result highlighting (in boldface and in red) follows the same rules as those described in Table II.
@@
Resistance to FL Non-i.i.d. We evaluate all baseline methods under varying degrees of non-i.i.d. data across three datasets, using a vanilla backdoor setting with a fixed spot-pattern trigger and a learning rate matching that of benign clients. As shown in Table II, Coward consistently achieves the best overall performance across all levels of heterogeneity, with TPRs above 95% and FPRs below 10.5% in all cases. Besides the overall conclusion, we draw several key observations: (1) Coward significantly reduces FPR across all settings compared to BackdoorIndicator, highlighting its advantage in mitigating OOD bias. A more detailed analysis is provided in Section V-E. (2) Proactive methods significantly outperform passive ones under severe data heterogeneity. Taking the most challenging non-i.i.d. setting ($\alpha=0.3$) as an example, aside from our method, BackdoorIndicator also outperforms all passive baselines, achieving TPRs of 93% on CIFAR-10 and 96% on EMNIST. This result supports the robustness of proactive defenses in heterogeneous environments. In contrast, passive methods exhibit significant performance degradation. For instance, the state-of-the-art passive method Flame sees its TPR drop from above 95% to below 10% on both CIFAR-100 and EMNIST, which aligns with our earlier analysis. (3) Some passive methods fail due to their unrealistic assumptions. For example, Rflbat performs poorly across all heterogeneity levels, likely because it relies only on the top two principal components of model parameters, losing critical information, especially under our challenging FL setup. Similarly, FoolsGold collapses on CIFAR-100, always judging all clients as benign (resulting in 0% TPR and 0% FPR). This failure stems from two factors: the high-dimensional 100-class prediction vector and its reliance on multi-attacker scenarios. These conditions lead to extremely low cosine similarity values, causing an all-as-benign judgment.
TABLE IV: Detection performance under multiple client attack on CIFAR-10 dataset. Result highlighting (in boldface and in red) follows the same rules in Table II.
@@
Resistance to Advanced Attacks. We evaluate all defense methods under more challenging client-side attacks. Specifically, across all attack methods, the malicious client adopts a more imperceptible noise-blended trigger and reduced local learning rate. These advanced configurations significantly increase the stealthiness of the attack. As shown in Table III, our method achieves the best overall performance, with proactive defenses showing strong robustness under advanced threats, while passive defenses suffer substantial degradation. This highlights the advantage of the proactive paradigm in handling sophisticated attacks. (1) Among proactive methods, Coward consistently achieves the best performance, maintaining close to 100% TPR and sub-10% FPR across all advanced attack types, including state-of-the-art strategies such as Neurotoxin and Chameleon. BackdoorIndicator also performs strongly, with TPRs consistently exceeding 85% in most settings; however, it suffers from relatively high FPRs, reflecting its vulnerability to distributional bias. (2) For passive methods, performance degradation is anticipated under such stealthy attacks, since the attacker introduces only subtle changes to the model, which reduces its suspiciousness. As an instance, MultiKrum, which selects representative updates based on Euclidean distance, becomes highly ineffective, with TPRs falling below 12.5% across all stealthy attack settings. Flame, which filters out updates with large angular deviations, also suffers notable degradation, with TPRs dropping below 30.5% on CIFAR-10 dataset. In contrast, DeepSight shows comparatively better resilience, which perform consistently with TPRs over 40% in most advanced attack cases. This improvement most probably stems from its joint use of both parameter and prediction discrepancies during its detection process.
@@
Robustness to Multiple Attackers. Table IV presents detection performance under multi-attacker scenarios. As shown in this table, proactive methods exhibit better overall performance across all three settings. This is expected as the detection mechanism of proactive approaches is independent of the number of attackers. Among them, our method achieves the best overall performance, consistently maintaining TPRs above 99% and FPRs below 15%. BackdoorIndicator also performs well but becomes less effective when the attacker proportion reaches 70%. In such cases, a temporary detection failure allows the increased number of attackers to rapidly reinforce the backdoor in the global model, resulting in a higher ASR. Moreover, the retained backdoor could be further maintained by the indicator’s own planting process due to the maintenance effect. Among passive baselines, FoolsGold performs best under Uniform and DBA attacks, achieving a TPR of 100%, benefiting from its assumption of highly similar, coordinated malicious updates. However, it fails in uncoordinated scenarios, where attackers pursue different targets. In such cases, the TPR drops below 40% as the method’s core assumption of shared attack objectives no longer holds. Outlier-based defenses such as MultiKrum and Flame become ineffective when attackers control 70% of clients, with TPRs dropping under 42% in all cases. These methods assume malicious updates are outliers, which fails when adversaries form the majority among all clients.
@@
TABLE V: Performance on different OOD planting sets. Result highlighting (in boldface and in red) follows the same rules as those described in Table II.
@@ V-C Ablation Study
Robustness to Choices of OOD Datasets. Table V presents detection performance across diverse OOD planting sets, including task-similar data, distinctly different data, and an extreme case with synthetic Gaussian noise. We observe that Coward consistently maintains low FPR across all settings, while BackdoorIndicator’s performance degrades as the OOD dataset diverges from the main task. Using the most challenging main task on CIFAR-100 dataset as example, Coward’s FPR remains below 6.3% across all planting sets. In contrast, BackdoorIndicator’s FPR increases sharply from 44.6% on CIFAR-10 to 60% with Gaussian noise. This confirms our earlier analysis of BackdoorIndicator’s increased vulnerability to OOD bias. Overall, these results demonstrate that Coward offers improved robustness to the choice of OOD data, enabling more flexible deployment in real-world scenarios.
Robustness to Choices of Triggers. Table VI summarizes detection performance under various combinations of trigger types. The attacker’s trigger varies in increasing stealthiness: from a visible pixel pattern, to an invisible blend pattern, and finally to an implicit semantic pattern naturally selected from the dataset. The watermark trigger also spans a range of types, including four visible patterns (diagonal, square, triangular, and noise), as well as two invisible variants: a mosaic pattern and the sample-specific WaNet pattern. We see that Coward performs consistently well across all settings, achieving a TPR above 96.5% and an FPR below 17.3%. These results demonstrate our method’s strong robustness to different triggers and the generalizability of collision effect.
TABLE VI: Detection performance with different combinations of server-side watermark trigger (S) and attackers’ client-side malicious trigger (A) on CIFAR-10.
@@
TABLE VII: Defense Performance of Coward under different watermark detection thresholds on CIFAR-10 dataset.
Threshold TPR FPR ASR BA
@@
Robustness to Choices of Thresholds. We further evaluate the impact of the detection threshold on the effectiveness of Coward. Since both attackers and benign clients may disrupt the watermark to some extent, the optimal threshold tends to fall within a relatively low range. Thus, we conduct threshold analysis within a mild range from 1% to 20%. As shown in Table VII, varying the detection threshold has minimal impact on TPR, which consistently remains above 98%. However, it progressively increases the FPR, rising from 1.0% to 13.9%. This increase is primarily due to client-side data heterogeneity. Certain benign clients, particularly those with severely skewed local distributions, may forget the watermark and exhibit significantly lower watermark accuracy than the majority. In contrast, malicious clients consistently show very low watermark accuracy. Empirically, we find that setting the threshold at 5% offers a reliable and practical guideline for a balanced TPR and FPR across diverse scenarios.
V-D Resistance to Potential Adaptive Attack
@@
Potential Adaptive Attack Strategy. We hereby consider adaptive attackers attempt to bypass our collision-based detection by injecting a similar watermark to preserve high watermark accuracy. Since the watermark target is chosen from a finite, known set, the attacker has a non-trivial chance of guessing it. To model this, we design a periodic guessing attack with switching period $T$, where the attacker updates the guessed label every $T$ rounds. Smaller $T$ increases the chance of early success but may cause unstable feedback, while larger $T$ offers more stable signals at the cost of later success.
@@
V-E Why is Our Method Effective?
To understand why our method works, we provide both overall and case-specific analyses on the CIFAR-10 dataset under the highly non-i.i.d. setting with $\alpha=0.3$.
Overall OOD Bias Analysis. We evaluate the overall bias level under vanilla FL, BackdoorIndicator, and our method, and further investigate its correlation with detection false positive rate. Specifically, the prediction bias is quantified by the standard deviation of class-wise prediction probabilities, since uniform predictions are ideal for OOD samples, and any bias toward a specific class results in non-uniformity, thereby increasing the variance. By aggregating the bias across all participating clients over all training epochs, we obtain the overall bias level throughout the training process. As results shown in Figure 7, it is evident that: (1) Our method yields the lowest bias distribution, validating the effectiveness of regulating the OOD mapping and highlighting the adverse effect of BackdoorIndicator’s random mapping strategy. (2) Even at high bias levels, our method achieves a lower FPR, demonstrating its robustness against OOD bias.

### v2

@@ 4.2 Watermark Interaction and Detection
5 Experiments
5.1 Main Settings
Datasets and Models. As a fundamental form of AI-generated categorical decisions, image classification has served as a standard benchmark for evaluating multimedia adversarial and forensic mechanisms. Following common practice in prior work, we conduct all experiments on three image benchmark datasets, EMNIST [cohen2017emnist], CIFAR-10, and CIFAR-100 [krizhevsky2009learning], using the ResNet-18 [he2016deep] architecture.
@@ 5.1 Main Settings
Federated Training Configurations. All experiments adopt the standard FedAvg setting [mcmahan2017fedavg], where we simulate 100 clients in total, with 10 randomly selected to participate in each training round. To simulate the inherent data heterogeneity across clients, we follow [li2020fedprox] to use Dirichlet sampling to partition the dataset. The Dirichlet concentration parameter $\alpha$ controls the label distribution skewness, where smaller values indicate more unbalanced distributions. Specifically, we use $\alpha=0.9$ by default to simulate mild skewness, and decrease it to 0.3 to model more severe distribution shifts. During local training, each benign client performs 2 epochs with a learning rate of 0.03, while the global training proceeds for 1200 rounds. For conciseness, we omit the less critical parameter details and follow [li2024backdoorindicator] for all settings not explicitly specified.
@@
Backdoor Manipulation Configurations. To ensure a comprehensive evaluation, we consider both manipulation scenarios of single-attacker and multi-attacker, each covering various attack strategies and backdoor types. (1) In the single-attacker scenario, we adopt a range of attack baselines, including the naive Vanilla training [gu2020badnets], the stealthy PGD attack [zitengsun01pgd], and the latest state-of-the-art methods Neurotoxin [Zhengming01Neurotoxin] and Chameleon [chameleon]. Regarding the backdoor type, we adopt three diverse types of triggers, ranging from the traditional visible trigger (BadNet [gu2020badnets]), to the invisible trigger (Blend [chen2017targeted]), and the natural feature-based trigger (Semantic [bagdasaryan01]), with increasing levels of imperceptibility. (2) In the multi-attacker scenario, we increase the ratio of attackers from 30% to 70%. To ensure comprehensive coverage, we include both aligned and divergent attack goals: Uniform [foolsgold] with homogeneous attackers, DBA [xie2019dba] with a shared target but distributed triggers, and NBA [nguyen2024nba] with non-cooperative attackers using distinct backdoors. Across all settings, poisoning uses a consistent learning rate of 0.03 by default, aligned with benign training, and reduces it to half for more challenging stealthy evaluation purposes. The backdoor target is class 0 by default. Similar to BackdoorIndicator, the malicious client launches the attack at the 1000th round, with poisoning action lasting for 200 rounds.
@@
Backdoor Detection Configurations. We evaluate six representative federated backdoor detection methods that span a range of strategies, from classical model-space analysis to more advanced and proactive techniques: We begin with traditional methods that rely on the statistical properties of model updates: (1) MultiKrum [multiKrum] detects manipulated models based on the Euclidean distance among client updates. (2) Foolsgold [foolsgold] identifies malicious clients by analyzing pairwise cosine similarity to detect overly aligned update directions. (3) Rflbat [wang2022rflbat] projects model updates into a PCA space and flags outliers as potential attackers. Then we further include more advanced approaches that incorporate prediction behavior or mitigation mechanisms: (4) Deepsight [rieger2022deepsight] evaluates similarity in model predictions and output neuron activations to reveal abnormal client behavior. (5) Flame [nguyen2022flame] combines detection based on model similarity together with mitigation techniques such as gradient clipping and noise injection. Finally, we include the pioneering proactive method that shifts the detection paradigm: (6) BackdoorIndicator [li2024backdoorindicator] injects OOD patterns into the global model and identifies manipulated models based on their reactions to these patterns.
@@
Configuration of Coward. Here we detail the key factors related to watermark injection and detection. To construct the OOD planting set, we use EMNIST (expanded to three channels) when the main task is CIFAR-10 or CIFAR-100, and grayscale CIFAR-10 when the main task is EMNIST. The same planting set is used for Indicator to ensure a fair comparison. By default, the planting set contains 1,000 samples. For watermark injection, we use 20% planting set samples for trigger-mapping, with a learning rate of 0.001 for 5 iterations. In our main experiments, we adopt WaNet [nguyen2021wanet] as the default watermark trigger. For detection, a fixed 5% threshold is used across all settings: clients with watermark accuracy below this are identified as malicious. The watermark target is set to 8 by default.
@@
5.2 Main Results
For a comprehensive evaluation, We evaluate all baseline detection methods from following three perspectives:
TABLE II: Detection performance against vanilla single-client manipulation attack under different non-i.i.d. settings. Boldface indicates the best result value (excluding the collapsed case with TPR/FPR=0/0), while red highlights metric-specific failures based on a 50% threshold ($i.e.$, values above 50% for FPR and ASR, and below 50% for TPR and BA).
@@ 5.2 Main Results
TABLE III: Detection performance against stealthy single-client manipulation attacks with advanced training strategies. Result highlighting (in boldface and in red) follows the same rules as those described in Table II.
@@
Resistance to FL Non-i.i.d. We evaluate all baseline methods under varying degrees of non-i.i.d. data across three datasets, using a vanilla backdoor setting with a fixed spot-pattern trigger and a learning rate matching that of benign clients. As shown in Table II, Coward consistently achieves the best overall performance across all levels of heterogeneity, with TPRs above 95% and FPRs below 10.5% in all cases. Besides the overall conclusion, we draw several key observations: (1) Coward significantly reduces FPR across all settings compared to BackdoorIndicator, highlighting its advantage in mitigating OOD bias. A more detailed analysis is provided in Section 5.5. (2) Proactive methods significantly outperform passive ones under severe data heterogeneity. Taking the most challenging non-i.i.d. setting ($\alpha=0.3$) as an example, aside from our method, BackdoorIndicator also outperforms all passive baselines, achieving TPRs of 93% on CIFAR-10 and 96% on EMNIST. This result supports the robustness of proactive detection in heterogeneous environments. In contrast, passive methods exhibit significant performance degradation. For instance, the state-of-the-art passive method Flame sees its TPR drop from above 95% to below 10% on both CIFAR-100 and EMNIST, which aligns with our earlier analysis. (3) Some passive methods fail due to their unrealistic assumptions. For example, Rflbat performs poorly across all heterogeneity levels, likely because it relies only on the top two principal components of model parameters, losing critical information, especially under our challenging FL setup. Similarly, FoolsGold collapses on CIFAR-100, always judging all clients as benign (resulting in 0% TPR and 0% FPR). This failure stems from two factors: the high-dimensional 100-class prediction vector and its reliance on multi-attacker scenarios. These conditions lead to extremely low cosine similarity values, causing an all-as-benign judgment which is unhelpful.
TABLE IV: Detection performance under multi-client backdoor manipulation on CIFAR-10 dataset. Result highlighting (in boldface and in red) follows the same rules in Table II.
@@
Resistance to Advanced Manipulations. We evaluate all baselines under more challenging client-side manipulations. Specifically, across all attack methods, the malicious client adopts a more imperceptible noise-blended trigger and reduced local learning rate. These advanced configurations significantly increase the stealthiness of the manipulation. As shown in Table III, our method achieves the best overall performance, with proactive detection showing strong robustness under advanced threats, while passive defenses suffer substantial degradation. This highlights the advantage of the proactive paradigm in handling sophisticated attacks. (1) Among proactive methods, Coward consistently achieves the best performance, maintaining close to 100% TPR and sub-10% FPR across all advanced attack types, including state-of-the-art strategies such as Neurotoxin and Chameleon. BackdoorIndicator also performs strongly, with TPRs consistently exceeding 85% in most settings; however, it suffers from relatively high FPRs, reflecting its vulnerability to distributional bias. (2) For passive methods, performance degradation is anticipated under such stealthy attacks, since the attacker introduces only subtle changes to the model, which reduces its suspiciousness. As an instance, MultiKrum, which selects representative updates based on Euclidean distance, becomes highly ineffective, with TPRs falling below 12.5% across all stealthy attack settings. Flame, which filters out updates with large angular deviations, also suffers notable degradation, with TPRs dropping below 30.5% on CIFAR-10 dataset. In contrast, DeepSight shows comparatively better resilience, which perform consistently with TPRs over 40% in most advanced attack cases. This improvement most probably stems from its joint use of both parameter and prediction discrepancies.
@@
Robustness to Multiple Manipulations. Table IV presents detection performance under multi-attacker scenarios. As shown in this table, proactive methods exhibit better overall performance across all three settings. This is expected as the detection mechanism of proactive approaches is independent of the number of attackers. Among them, our method achieves the best overall performance, consistently maintaining TPRs above 99% and FPRs below 15%. BackdoorIndicator also performs well but becomes less effective when the attacker proportion reaches 70%. In such cases, a temporary detection failure allows the increased number of attackers to rapidly reinforce the backdoor in the global model, resulting in a higher ASR. Moreover, the retained backdoor could be further maintained by the indicator’s own planting process due to the maintenance effect. Among passive baselines, FoolsGold performs best under Uniform and DBA attacks, achieving a TPR of 100%, benefiting from its assumption of highly similar, coordinated malicious updates. However, it fails in uncoordinated scenarios, where attackers pursue different targets. In such cases, the TPR drops below 40% as the method’s core assumption of shared attack objectives no longer holds. Outlier-based defenses such as MultiKrum and Flame become ineffective when attackers control 70% of clients, with TPRs dropping under 42% in all cases. These methods assume malicious updates are outliers, which fails when adversaries form the majority among all clients.
@@
TABLE V: Performance on different OOD planting sets. Result highlighting (in boldface and in red) follows the same rules in Table II.
@@ 5.3 Ablation Study
Robustness to Choices of OOD Datasets. Table V presents detection performance across diverse OOD planting sets, including task-similar data, distinctly different data, and an extreme case with synthetic Gaussian noise. We observe that Coward consistently maintains low FPR across all settings, while Indicator’s performance degrades as the OOD dataset diverges from the main task. Using the most challenging main task on CIFAR-100 dataset as example, Coward’s FPR remains below 6.3% across all planting sets. In contrast, Indicator’s FPR increases sharply from 44.6% on CIFAR-10 to 60% with Gaussian noise. This confirms our earlier analysis of Indicator’s increased vulnerability to OOD bias. Overall, these results demonstrate that Coward offers improved robustness to the choice of OOD data, enabling more flexible deployment in real-world scenarios.
Robustness to Choices of Triggers. Table VI summarizes detection performance under various combinations of trigger types. The attacker’s trigger varies in increasing stealthiness: from a visible pixel pattern, to an invisible blend pattern, and finally to an implicit semantic pattern naturally selected from the dataset. The watermark trigger also spans a range of types, including four visible patterns (diagonal, square, triangular, and noise), as well as two invisible variants: a mosaic pattern and the sample-specific WaNet pattern. We see that Coward performs consistently well across all settings, achieving a TPR above 96.5% and an FPR below 17.3%. These results demonstrate our method’s strong robustness to different triggers and the generalizability of the multi-backdoor collision effect.
TABLE VI: Detection performance with different combinations of watermark trigger (S) and attackers’ client-side malicious trigger (A) on CIFAR-10.
@@
TABLE VII: Detection performance of Coward under different watermark-retention thresholds on the CIFAR-10 dataset.
[…]
@@
Robustness to Choices of Detection Thresholds. We further evaluate the impact of the detection threshold on the effectiveness of Coward. Since both attackers and benign clients may disrupt the watermark to some extent, the optimal threshold tends to fall within a relatively low range. Thus, we conduct threshold analysis within a mild range from 1% to 20%. As shown in Table VII, varying the detection threshold has minimal impact on TPR, which consistently remains above 98%. However, it progressively increases the FPR, rising from 1.0% to 13.9%. This increase is primarily due to client-side data heterogeneity. Certain benign clients, particularly those with severely skewed local distributions, may forget the watermark and exhibit significantly lower watermark accuracy than the majority. In contrast, malicious clients consistently show very low watermark accuracy. Empirically, we find that setting the threshold at 5% offers a reliable and practical guideline for a balanced TPR and FPR across diverse scenarios.
5.4 Resistance to Potential Adaptive Attack
@@
Potential Adaptive Attack Strategy. We consider adaptive attackers attempt to bypass our collision-based detection by injecting a similar watermark to preserve high watermark accuracy. Since the watermark target is chosen from a finite, known set, the attacker has a non-trivial chance of guessing it. To model this, we design a periodic guessing attack with switching period $T$, where the attacker updates the guessed label every $T$ rounds. Smaller $T$ increases the chance of early success but may cause unstable feedback, while larger $T$ offers more stable signals at the cost of later success.
@@
5.5 Why is Our Method Effective?
To understand why our method works, we provide both overall and case-specific analyses on the CIFAR-10 dataset under a highly non-IID distribution with $\alpha=0.3$.
Overall OOD Bias Analysis. We evaluate the overall bias level under vanilla FL, BackdoorIndicator, and our method, and further investigate its correlation with detection false positive rate. Specifically, the prediction bias is quantified by the standard deviation of class-wise prediction probabilities, since uniform predictions are ideal for OOD samples, and any bias toward a specific class results in non-uniformity, thereby increasing the variance. By aggregating the bias across all participating clients over all training epochs, we obtain the overall bias level throughout the training process. As results shown in Figure 7, it is evident that: (1) Our method yields the lowest bias distribution, validating the effectiveness of regulating the OOD mapping and highlighting the adverse effect of BackdoorIndicator’s random mapping strategy. (2) Even at comparable bias levels, our method achieves a lower FPR, demonstrating its robustness against OOD bias.